OpenStudy (anonymous):

I know how to do this, I jus need the equation set up. I = RPT I think you deposited $600 dollars in a savings account that earns a simple interest rate of 3.5% per year. How many years do you need to keep the money in the account to earn $210 in interest ?

jimthompson5910 (jim_thompson5910):

I've seen it as I = PRT, but I = RPT works too

jimthompson5910 (jim_thompson5910):

In this case, I = 210 P = 600 R = 0.035 T = unknown (you're solving for this)

OpenStudy (anonymous):

Oh that's probably how it is then. I'm jus confused because the rate is the percent and there isn't another. I tried I = 3.5% ($600)$210 but it didn't look right

jimthompson5910 (jim_thompson5910):

3.5% = 3.5/100 = 0.035

jimthompson5910 (jim_thompson5910):

it should be I = PRT 210 = 600*0.035*T solve for T

OpenStudy (anonymous):

Ok so how would the equation look

jimthompson5910 (jim_thompson5910):

it looks like this 210 = 600*0.035*T

jimthompson5910 (jim_thompson5910):

from there you need to solve for T

OpenStudy (anonymous):

Ok thanks. so jus to make sure, I would multiply 600(3.5) then divide the answer by 210 ?

jimthompson5910 (jim_thompson5910):

600 times 0.035 = ???

OpenStudy (anonymous):

I got 2100 so that's not right

jimthompson5910 (jim_thompson5910):

no that's way too high

jimthompson5910 (jim_thompson5910):

you have the decimal in the wrong spot

OpenStudy (anonymous):

No. it's 600(3.5) that's how it's written on the paper so that's the way I'm solving it.

jimthompson5910 (jim_thompson5910):

ok well you need to divide by 100 because 3.5% = 3.5/100

OpenStudy (anonymous):

Actually I have no idea how you're getting0.035

jimthompson5910 (jim_thompson5910):

3.5% = 3.5/100 = 0.035 therefore 3.5% = 0.035

jimthompson5910 (jim_thompson5910):

because that's what "percent" means

OpenStudy (anonymous):

Ok well I'm not trying to solve 3.5 I'm trying to solve the whole thing.

jimthompson5910 (jim_thompson5910):

percent = per hundred example 25% = 25/100 = 0.25

jimthompson5910 (jim_thompson5910):

yes but you don't use 3.5, you use 0.035 because you need to convert from a percent to a decimal

OpenStudy (anonymous):

Ok I see where I/U went wrong on that. u move thel the other way and jus make it 35

jimthompson5910 (jim_thompson5910):

no you move the decimal 2 spots to the left another example 72% = 0.72

jimthompson5910 (jim_thompson5910):

so saying "72% of 25" is the same as saying 0.72*25

OpenStudy (anonymous):

I don't need an example. I need the proper equation so I can solve it.

jimthompson5910 (jim_thompson5910):

I gave it to you 210 = 600*0.035*T

jimthompson5910 (jim_thompson5910):

you don't need to be rude